OpenStudy (anonymous):

how do you expand: log base 10 square root of x^3

OpenStudy (rulnick):

1.5 log_10 (x)

OpenStudy (anonymous):

\[\log_{10} \sqrt{x^3}\] How do you expand that?

OpenStudy (rulnick):

It's the same as log base 10 of x to the 3/2 power, which then equals what I wrote above.

OpenStudy (anonymous):

are you sure? can you explain?

OpenStudy (rulnick):

log a^b = b log a

OpenStudy (rulnick):

for you, b is 3/2

OpenStudy (rulnick):

because square root of x^3 is x^(3/2)

OpenStudy (accessdenied):

\[ \sqrt{x} = x^{1/2} \]

OpenStudy (anonymous):

so it would be 3/2 long base 10 x?

OpenStudy (rulnick):

Yup.

OpenStudy (anonymous):

okay thanks